### v1

# FLARE: Agile Flights for Quadrotor Cable-Suspended Payload System via Reinforcement Learning

Abstract
Agile flight for the quadrotor cable-suspended payload system is a formidable challenge due to its underactuated, highly nonlinear, and hybrid dynamics. Traditional optimization-based methods often struggle with high computational costs and the complexities of cable mode transitions, limiting their real-time applicability and maneuverability exploitation. In this letter, we present FLARE, a reinforcement learning (RL) framework that directly learns agile navigation policy from high-fidelity simulation. Our method is validated across three designed challenging scenarios, notably outperforming a state-of-the-art optimization-based approach by a 3x speedup during gate traversal maneuvers. Furthermore, the learned policies achieve successful zero-shot sim-to-real transfer, demonstrating remarkable agility and safety in real-world experiments, running in real time on an onboard computer. A supplementary video is available at [video [^2] https://www.youtube.com/playlist?list=PLbEQeDMEVpqxrKderAjZ3R-KmNCdQaXvUK].
{strip}[Uncaptioned image]
Figure 1: Autonomous Agile Flights of a Quadrotor Cable-suspended Payload System via Reinforcement Learning Policy. The system demonstrates its capabilities in three challenging real-world scenarios: (a) agile waypoint passing, (b) payload targeting, and (c1, c2) one-time and sequential gate traversal.
I INTRODUCTION
Over the past few decades, agile flight technology for quadrotors has advanced rapidly. More recently, there appears to be growing research interest in agile flight of quadrotor-based aerial robotic platforms. Among these platforms, cable-suspended payload systems have attracted considerable attention, with a burgeoning field of research exploring their promising applications [1], [2], [3].
Navigating the suspended payload system has always been a challenging problem, let alone achieving agile flights. The inherent strong nonlinearity and underactuation of the system impose fundamental difficulties in planning and control. Moreover, the system exhibits hybrid dynamic characteristics, where the taut or slack mode of the cable significantly alters the system dynamics, presenting additional challenges. Driven by this concern, some early studies [4], [5] focus primarily on minimizing cable oscillations to ensure safety at the expense of flexibility and agility.
To solve this problem, various approaches based on optimization are proposed, aiming to further exploit the system’s maneuverability for more agile flights. These approaches converge on the idea of planning motions by solving an optimal control problem (OCP) in a model-based manner. Within these approaches, the cable swing dynamics is actively leveraged to facilitate agile flights of the suspended payload system [6], even achieving collision-free agile flights in cluttered environments [7], [8]. Despite their impressive results, these works are predicated exclusively on the taut-cable paradigm, without considering the slack mode. To address this limitation, a unified system model is introduced that incorporates complementarity constraints to describe mode transitions [9]. Consequently, some works exploit the hybrid modes of this system to generate or track agile trajectories involving automatic mode switching [10], [11]. More recently, Wang et al. achieve flights with automatic mode-switching through a systematic planning and control framework, which is remarkably validated in a gate-traversal scenario [12]. However, the corresponding high computational costs prevent real-time trajectory planning in these works [9, 10, 12].
Notably, Reinforcement Learning (RL) method offers a promising alternative in a data-based manner, with extensively growing applications in quadrotor agile flights in recent years [13], [14], [15]. Indeed, applying RL methods to achieve agile flights for the suspended payload system offers three principal advantages. First, the model-free nature of RL eliminates reliance on gradients of the dynamic model, thereby avoiding the significant difficulties encountered by traditional gradient-based optimization methods when handling nonsmooth problems. Second, the stochastic exploration mechanisms embedded in most RL algorithms grant them the ability to escape local minima, crucial for solving non-convex optimization problems. Third, the trained policies are typically computationally lightweight, permitting their onboard deployment for real-time planning.
Meanwhile, RL methods have been preliminarily explored for the navigation of the suspended payload system. Pioneering RL implementations prioritize the swing minimization [16], [17], achieving safety guarantees while compromising on agility. Another work augments the RL framework with meta-learning technique, facilitating rapid mid-flight adaptation to unknown physical properties [18], achieving robust performance across varying payload weights and cable lengths. However, the application of RL methods to achieve truly agile flights of the suspended payload system still remains unexplored.
Hence, in this letter, we tackle the challenge of agile flight for the suspended payload system by formulating it within the RL framework. We employ a model-free RL method in a high-fidelity, high-performance simulation platform. The trained policy network establishes a direct state-to-control mapping, circumventing the need for separate planning and control modules. Additionally, the lightweight network enables high-frequency inference solely with onboard resources, which is challenging for computationally intensive optimization-based approaches. Overall, the main contributions of this letter are summarized as follows:
  1. 1.
    We propose an RL-based navigation framework for the agile flight of the quadrotor cable-suspended payload system, featuring reward functions tailored to the characteristics of the system.
  2. 2.
    We design three representative application scenarios and instantiate the RL problem to demonstrate the method’s promising capabilities.
  3. 3.
    We validate our proposed method and its zero-shot sim-to-real transfer capability through extensive experiments conducted in both simulation and the real world. Our approach demonstrates remarkable agility and safety with high-performance, real-time execution.
II METHODOLOGY
Figure 2: Overview of our proposed approach. We design three representative application scenarios focused on motion planning of cable-suspended payload system, each with a unique goal. In each scenario, a control policy is learned with customized observations and reward setting in Genesis simulator [19].
This section details our agile but safe motion planning approach for the quadrotor suspended payload system. We first present the system’s complex hybrid dynamics, the challenges of which motivate our model-free RL approach. We then formulate the motion planning problem as a Markov Decision Process and design three representative application scenarios: agile waypoint passing, payload targeting, and gate traversal. Finally, we describe more training details, including a domain randomization strategy to enhance the policy’s real-world generalization.
II-A Suspended-payload system dynamics
The quadrotor suspended system exhibits hybrid dynamics characterized by two modes depending on whether the cable is taut or slack.
In the taut mode, the system dynamics is governed by:
$(m_{q}+m_{p})(\boldsymbol{\ddot{x}_{p}}-\boldsymbol{g})=\left(\boldsymbol{\rho}\cdot\mathbf{R}\boldsymbol{T}-m_{q}l(\boldsymbol{\dot{\rho}}\cdot\boldsymbol{\dot{\rho}})\right)\boldsymbol{\rho}$ (1)
$\boldsymbol{x_{q}}=\boldsymbol{x_{p}}-l\boldsymbol{\rho},~~\dot{\mathbf{R}}=\mathbf{R}\boldsymbol{\hat{\omega}}$
$m_{q}l\boldsymbol{\ddot{\rho}}=-m_{q}l(\boldsymbol{\dot{\rho}}\cdot\boldsymbol{\dot{\rho}})\boldsymbol{\rho}+\boldsymbol{\rho}\times(\boldsymbol{\rho}\times\mathbf{R}\boldsymbol{T})$
$I_{q}\boldsymbol{\dot{\omega}}=\boldsymbol{M}-\boldsymbol{\omega}\times I_{q}\boldsymbol{\omega}$
where $\mathbf{T}=[0,0,T]^{T}$ represents the thrust vector, $\mathbf{R}$ denotes the rotation matrix, $\mathbf{g}$ is the gravity vector, $l$ is the cable length, $\boldsymbol{\rho}$ denotes the cable direction unit vector from quadrotor to payload. The variables $\mathbf{x}_{q}$, $\boldsymbol{\omega}$, $m_{q}$, $\mathbf{I}_{Q},\boldsymbol{M}$ represent the postion, body angular rates, mass, the inertia matrix and the moment vector of quadrotor, and $\mathbf{x}_{p}$, $m_{p}$ represent the postion and mass of payload, respectively.
In the slack mode (zero cable tension), the quadrotor resumes to nominal quadrotor dynamics, while the payload obeys free-fall dynamics.
These complex hybrid dynamics make optimization-based methods computationally expensive. Therefore, this work employs a model-free RL approach to achieve real-time navigation.
II-B Problem formulation
We formulate the motion planning problems of the suspended payload system as a general infinite-horizon Markov Decision Process (MDP), defined as $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\gamma)$. Here, $\mathcal{S}$ denotes the state space, $\mathcal{A}$ the action space, $\mathcal{P}$ the transition probability, $\mathcal{R}$ the reward function, and $\gamma$ the discount factor. Within this MDP framework, we aim to optimize the parameters $\psi$ of a policy network $\pi_{\psi}$ to maximize the expected discounted return:
$\pi_{\psi}^{*}=\operatorname*{arg\,max}_{\pi}~\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}r(t)\right]$ (2)
where $\gamma\in[0,1)$ is the discount factor, and $r(t)$ represents the immediate reward at time step $t$. Within this framework, we develop three distinct application scenarios focused on suspended system motion planning, each presenting unique challenges. Following this, we define the observations, actions, and reward functions, as the general framework illustrated in Fig. 2.
II-B1 Observations and Actions
For all three scenarios, the observation vector $\mathbf{o}$ consists of two parts, defined as follows:
$\mathbf{o}=[\mathbf{o}_{\textit{general}},\mathbf{o}_{\textit{specific}}]^{T},\\$ (3)
where $\mathbf{o}_{\textit{general}}$ represents the general observations shared across all three scenarios, while $\mathbf{o}_{\textit{specific}}$ consists of scenario-specific observations that will be detailed in section II-C.
Specifically, the general observation component, $\mathbf{o}_{\textit{general}}=[\mathbf{o}_{q},\mathbf{o}_{p}]^{T}$, comprises the observations of quadrotor states $\mathbf{o}_{q}$ and the payload states $\mathbf{o}_{p}$. The quadrotor state observation, $\mathbf{o}_{q}=[\mathbf{v}_{q},\operatorname{vec}{\left(\mathbf{R}(\mathbf{q})\right)}]\in\mathbb{R}^{12}$, includes quadrotor velocity $\mathbf{v}_{q}$ and flattened rotation matrix $\operatorname{vec}{\left(\mathbf{R}(\mathbf{q})\right)}$ derived from quadrotor quaternion $\mathbf{q}$. The payload observation $\mathbf{o}_{p}=[\phi,\theta]^{\top}\in\mathbb{R}^{2}$ encodes its deviation angles from the negative z-axis $\mathbf{z}_{B}$ of the quadrotor body frame $\mathcal{B}$, (Fig. LABEL:cable-angle), where:
$\phi$ $=\mathrm{atan2}(y_{b},-z_{b}),~~\theta=\mathrm{atan2}(x_{b},-z_{b}),$ (4)
with $x_{p}^{b},y_{p}^{b},z_{p}^{b}]^{T}$ being the positional coordinates of the payload in body frame $\mathcal{B}$. These components of the general observation $\mathbf{o}_{\textit{general}}$ provide the policy with essential information about the system’s current states.
The policy network $\pi_{\psi}$ generates normalized actions $\mathbf{a}=[\tilde{T}_{\text{cmd}},\tilde{\boldsymbol{\omega}}_{\text{cmd}}]$, which are transformed to physical commands through:
$T_{\text{cmd}}=\frac{\tilde{T}_{\text{cmd}}+1}{2}T_{\text{max}},\quad\boldsymbol{\omega}_{\text{cmd}}=\tilde{\boldsymbol{\omega}}_{\text{cmd}}\boldsymbol{\omega}_{\text{max}},$ (5)
where $T_{\text{max}}$ and $\boldsymbol{\omega}_{\text{max}}$ represent the thrust-to-weight ratio and maximum body rate, respectively.
II-B2 Reward Function
Analogous to the observation vector, the reward function in all three scenarios is also composed of two parts, defined as follows:
${r}={r}_{\text{general}}+{r}_{\text{specific}},$ (6)
where ${r}_{\text{general}}$ denotes the general reward component common to all three scenarios, while ${r}_{\text{specific}}$ is scenario-specific and will be introduced in section II-C.
Specifically, the general reward component is designed as ${r}_{\textit{general}}={r}_{\textit{safe}}+{r}_{\textit{crash}}+{r}_{\textit{smooth}}$, where we difine the sub-rewards as follows.
The safety reward ${r}_{\text{safe}}$ is designed to mitigate the risk of catastrophic failure during agile flights. When the payload swings to a high altitude, the cable potentially entangles with the propellers. To this end, ${r}_{\text{safe}}$ imposes penalties when the payload’s deviation angle $\phi$ or $\theta$ (Fig. LABEL:cable-angle) exceeds the threshold $\phi_{max}$ (Fig. LABEL:unsafe_example), defined as:
[…]
The crash reward ${r}_{\text{crash}}$ penalizes boundary violations with episodic termination strategies. Specifically, the crash reward activates exclusively when the quadrotor or the payload violates the spatially invalid workspace $\mathcal{W}$, defined as:
[…]
The smoothness reward ${r}_{\text{smooth}}$ penalizes dynamically infeasible commands through limiting the abrupt variation of consecutive actions $\mathbf{a}$, defined as:
${r}_{\text{smooth}}=-\lambda_{1}{\left\|\mathbf{a}_{t-1}-\mathbf{a}_{t}\right\|}.$ (9)
II-C Scenario customization
In this section, we detail the scenario-specific components of the observation vector and reward function for each of the three scenarios, thereby completing their respective RL problem formulations. The training pipeline and the descriptions of these scenarios are depicted in Fig. 2.
II-C1 Agile waypoint passing
The goal of this scenario is to establish a baseline for agile navigation, testing the policy’s core capability for high-speed, safe traversal through a sequence of predefined waypoints.
To this end, the observation vector in the waypoint-passing scenario is defined as $\mathbf{o}^{\textit{WP}}=[\mathbf{o}_{\textit{general}},\mathbf{o}^{\textit{WP}}_{\textit{specific}}]^{T}$, where $\mathbf{o}^{\textit{WP}}_{\textit{specific}}=\mathbf{o}_{w}^{\textit{WP}}$ denotes the observation of waypoints. The waypoint observation is composed of the relative position vectors, $\Delta\mathbf{x}_{q_{k}},k\in\{1,...,W\}$, from the drone to the next $W$ waypoints. We set $W=2$ to take two future waypoints into consideration, resulting in the waypoint observation vector $\mathbf{o}_{w}^{\textit{WP}}=[\Delta\mathbf{x}_{q_{1}},\Delta\mathbf{x}_{q_{2}}]^{\top}\in\mathbb{R}^{6}$.
Furthermore, we design the reward function in this scenario as:
${r}^{\textit{WP}}={r}_{\text{general}}+\lambda_{2}r^{\textit{WP}}_{\text{specific}},$ (10)
${r}^{\textit{WP}}_{\text{specific}}={r}^{\textit{WP}}_{\text{target}},$
where the target reward ${r}^{\textit{WP}}_{\text{target}}$ quantifies the spatial progress toward the next waypoint, incentivizing the system to reach the target in minimum time, a key characteristic of agile flight, defined as:
$r^{\textit{WP}}_{\text{target}}=\|\Delta\mathbf{x}_{q_{t-1}}\|^{2}-\|\Delta\mathbf{x}_{q_{t}}\|^{2}.$ (11)
II-C2 Payload targeting
In this scenario, we redefine the operational paradigm from direct quadrotor navigation to payload targeting. The primary objective shifts from guiding the drone itself to directing the payload to its destination, which requires actively leveraging the system’s complex swing dynamics.
Similarly to scenario 1, the observation vector in payload-targeting scenario is defined as $\mathbf{o}^{\textit{PT}}=[\mathbf{o}_{\textit{general}},\mathbf{o}^{\textit{PT}}_{\textit{specific}}]^{T}$, where $\mathbf{o}^{\textit{PT}}_{\textit{specific}}=\mathbf{o}_{w}^{\textit{PT}}$ denotes the observation of waypoints for the payload. Since only one waypoint is given as the payload target, we define the waypoint observation as $\mathbf{o}_{w}^{\textit{PT}}=[\Delta\mathbf{x}_{q},\Delta\mathbf{x}_{p}]^{T}$, where $\Delta\mathbf{x}_{q}$, $\Delta\mathbf{x}_{p}$ are defined as the drone’s and the payload’s relative position to the target, respectively.
To manipulate the payload to reach a designated waypoint, we design the reward function as:
${r}^{\textit{PT}}={r}_{\text{general}}+\lambda_{2}r^{\textit{PT}}_{\text{specific}},$ (12)
${r}^{\textit{PT}}_{\text{specific}}={r}^{\textit{PT}}_{\text{target}},$
where the target reward ${r}^{\textit{PT}}_{\text{target}}$ in this scenario quantifies the spatial progress of the payload, instead of the quadrotor, towards the designated waypoint, defined as:
$r^{\textit{PT}}_{\text{target}}=\|\Delta\mathbf{x}_{p_{t-1}}\|^{2}-\|\Delta\mathbf{x}_{p_{t}}\|^{2}.$ (13)
II-C3 Gate traversal
This scenario assesses the policy’s capability for highly aggressive maneuvers under tight spatial constraints, where collision-free passage requires exploiting the full system dynamics. In this scenario, we aim to achieve a collision-free navigation through a narrow gate, with a given final waypoint position $\mathbf{p}_{w}$ for the quadrotor and a known gate position $\mathbf{p}_{g}$.
The observation vector in the gate-traversal scenario is also defined as $\mathbf{o}^{\textit{GT}}=[\mathbf{o}_{\textit{general}},\mathbf{o}^{\textit{GT}}_{\textit{specific}}]^{T}$, where $\mathbf{o}^{\textit{GT}}_{\textit{specific}}=[\mathbf{o}_{w}^{\textit{GT}},\mathbf{o}_{g}^{\textit{GT}}]^{T}$ encodes the observation of waypoints and the gate position, respectively. The waypoint observation $\mathbf{o}_{w}^{\textit{GT}}$ shares the identical definiton with $\mathbf{o}_{w}^{\textit{WP}}$. Additionally, we define the gate observation $\mathbf{o}_{g}^{\textit{GT}}$ as the drone’s relative position to the gate center $\mathbf{x}_{g}$, providing information about the gate’s position.
To achieve collision-free navigation through a narrow gate, we design the reward function as:
${r}^{\textit{GT}}={r}_{\text{general}}+\lambda_{2}r^{\textit{GT}}_{\text{specific}},$ (14)
${r}^{\textit{GT}}_{\text{specific}}={r}^{\textit{GT}}_{\text{target}}+\lambda_{3}{r}^{\textit{GT}}_{\text{gate}},$
where the target reward ${r}^{\textit{GT}}_{\text{target}}$ shares the the identical definition with ${r}^{\textit{WP}}_{\text{target}}$ in Equation 11, and the gate reward ${r}^{\textit{GT}}_{\text{gate}}$ incentivizes the drone to approach and traverse the gate with a two-stage design strategy, defined as:
[…]
$r_{guide}=\frac{(\mathbf{p}_{w}-\mathbf{p}_{g})}{\|\mathbf{p}_{w}-\mathbf{p}_{g}\|}\cdot\mathbf{v}_{q},$
where $r_{guide}$ incentives the magnitude of the drone’s velocity projected onto the desired direction, which is defined as a vector from the gate position to the waypoint position.
II-D Policy training
To ensure stable and efficient policy training in light of the varying scales of the input features, all observations are normalized prior to network processing: the quadrotor relative positions $\Delta\mathbf{x}_{q}\leftarrow\Delta\mathbf{x}_{q}/\mathbf{k}_{q}$, the quadrotor velocity $\mathbf{v}_{q}\leftarrow\mathbf{v}_{q}/\mathbf{k}_{v}$, the payload relative positions $\Delta\mathbf{x}_{p}\leftarrow\Delta\mathbf{x}_{p}/\mathbf{k}_{p}$, the gate relative positions $\mathbf{o}_{g}\leftarrow\mathbf{o}_{g}/\mathbf{k}_{g}$ and the deviation angles of the payload $\mathbf{o}_{p}\leftarrow\mathbf{o}_{p}/\mathbf{k}_{\phi}$, where the division is conducted element-wise, and $\mathbf{k}_{q}$, $\mathbf{k}_{v}$, $\mathbf{k}_{p}$, $\mathbf{k}_{g}$ and $\mathbf{k}_{\phi}$ are normalization constants listed in Table I.
TABLE I: Simulation and Reward Parameters
  Parameter Value Parameter Value
$T_{\text{max}}$ $3.5$ $\boldsymbol{\omega}_{\text{max}}$ [rad/s] $[15,15,5]$
$\mathbf{k}_{q},\mathbf{k}_{p}$ [m] $[5,5,1]$ $\mathbf{k}_{v}$ [m/s] $[10,10,3]$
$\mathbf{k}_{g}$ [m] $[3,3,1]$ $\mathbf{k}_{\phi}$ [rad] $[1.5,1.5]$
$r_{\text{bound}}$ $10$ $r_{\text{excess}}$ 3
$r_{\text{arrival}}$ 2020 $\lambda_{1}$ $1\times 10^{\text{-}4}$
$\lambda_{2}$ $10$ $\lambda_{3}$ $5\times 10^{-3}$
With the overall framework illustrated in Fig. 2, our policy is trained using Proximal Policy Optimization (PPO) [20]. The policy network, implemented as an MLP, consists of two hidden layers with 128 units each. A final projection layer then maps the resulting features to the normalized action space via a Tanh activation function. To better satisfy the Markov property, the network’s input observation is augmented with the action from the previous timestep.
The complex characteristics of the system typically incur a significant time cost. To address this, we leverage Genesis [19], a state-of-the-art, GPU-vectorized simulator that provides high-performance quadrotor simulation. Within this simulator, we model the system using the Unified Robot Description Format (URDF). The cable is specifically simulated as a serial chain of interconnected rigid links, an approach similar to that in [12]. All simulations are executed with a timestep of 0.01 s. This setup allows the policies for Scenarios 1, 2, and 3 to converge after approximately 0.75, 0.90, and 1.85 million timesteps, which corresponds to only 0.45, 0.52, and 2.42 hours of training time, respectively.
Moreover, the system’s complex dynamics may also introduce unmodeled factors that result in non-negligible sim-to-real gaps, such as unpredictable payload swings induced by external air disturbances and imperfect drone hovering. Therefore, we adopt a domain randomization strategy by incorporating random minor variations in the payload’s initial deviation angle, to ensure the distribution of training data encompasses real-world conditions.
III SIMULATION RESULTS AND ANALYSIS
This section presents the simulation results and further analysis of our approach. In scenario 1, we evaluate the performance of our trained policy across three tracks, utilizing a near-time-optimal policy for a single drone (without a suspended payload) on the same tracks as a reference. In scenario 2, we validate the method’s ability to navigate the payload to a given target and demonstrate that the policy effectively exploits the cable’s natural swing characteristics. Finally, in scenario 3, we benchmark the performance of our method against a state-of-the-art work [12] executing identical one-time gate traversals. All training and simulation experiments are performed on a workstation with an AMD Ryzen R9-9950 CPU and NVIDIA RTX 5090 D GPU.
III-A Agile waypoint passing
In this scenario, we validate the policy through numerous trials on tracks with varying complexity, where the drone is required to pass a given sequence of waypoints within a predefined radius.
We first rule out a naive approach: agile flights for the suspended payload system can be achieved simply by treating the payload as an external disturbance and relying on the inherent robustness of the RL method. To verify this, we additionally establish a reference policy only incorporating the drone dynamics, similar to the methods in [13], [15]. When applied to scenario 1, the policy trained solely on the drone’s dynamics is unable to complete the waypoint-passing task, often resulting in catastrophic failure (e.g. a crash), as illustrated in Fig. LABEL:sim_task1_a. In contrast, our policy, which accounts for the complete dynamics of the coupled system, successfully navigates all waypoints.
TABLE II: Performance Metrics Across Different Simulation Tracks in Scenario 1 (0.5m Threshold)
Configuration Performance Metrics
Track Dynamic Max Vel. Avg Vel. Time
Model (m/s) (m/s) (s)
Zigzag with payload $5.59$ $3.39$ $2.20$
single drone $6.47$ $3.75$ $1.68$
Eight - 3 laps with payload $5.64$ $3.38$ $11.96$
single drone $5.87$ $3.66$ $10.05$
Heart with payload $5.96$ $4.04$ $8.74$
single drone $7.11$ $4.62$ $7.42$
Furthermore, we conduct a focused analysis of the system’s agility in this scenario. We navigate the suspended system through three tracks and demonstrate the trajectories in Fig. LABEL:sim_task1_b. We also provide the performance of the aforementioned reference policy when it operates a single drone (without a payload) on the same waypoint-passing tracks. The key metrics summarized in Table II underscore the high agility of our method. While ensuring successful and safe task completion, our policy achieves an average maximum velocity of 88.38% and an average velocity of 89.86% relative to the reference policy, which is a near-time-optimal policy for single-drone navigation using the identical quadrotor configuration.
III-B Payload targeting
In this scenario, we validate the effectiveness of navigating the payload to reach a given target.
As shown in Fig. LABEL:genesis_task2 and Fig. LABEL:task2_method2, our method successfully navigates to a sequence of 10 payload targets, forming a double-loop, star-shaped trajectory. Additionally, we compare the velocity performance and completion time with those of scenario 1, where the identical targets are treated as quadrotor waypoints, as summarized in Table III.
TABLE III: Performance Metrics Comparison on 10 Identical Targets (0.2m Threshold)
Scenario Max Vel. Avg Vel. Time
(m/s) (m/s) (s)
1 $4.05$ $2.55$ $8.57$
2 $3.33$ $1.78$ $8.15$
The aforementioned comparison of the policies from scenario 1 and scenario 2 reveals that the latter learns to exploit the natural swing characteristics of the cable. This learned behavior allows the payload to reach targets by harnessing its momentum, resulting in a more efficient flight with a shorter travel distance and duration.
III-C Gate traversal
TABLE IV: Performance Metrics Comparison Over 8 Simulation Tracks: Maximum Velocity, Traversal Time, Computational Time, Quadrotor Traversal Deviation, Payload Traversal Deviation, and Success Rate.
Method Average Performance Metrics (Mean ± Std. Dev.) Success Rate(%)
Max Vel. (m/s) Time (s) Comp. Time (s) Quad. Dev. (m) Payload Dev. (m)
Impactor 1.285 ± 0.087 2.082 ± 0.099 4.308 ± 1.418 0.181 ± 0.023 0.169 ± 0.055 100
Ours 3.673 ± 0.446 0.771 ± 0.082 $\leq$~0.01 0.169 ± 0.041 0.144 ± 0.048 100
In this scenario, we evaluate the performance of our method with a challenging gate traversal maneuver. The objective is to guide both the quadrotor and suspended payload through a circular gate with a radius of 0.3 m, ensuring collision-free passage.
Figure 6: System motions of gate traversal in scenario 3.
We validate our method’s ability by performing both one-time and sequential gate traversal, as shown in Fig. 6. For performance evaluation, we benchmark our approach against Impactor [12] in the one-time traversal setting. This comparison is conducted across 8 trials, where the final waypoint is positioned within the range of [-2.0, 2.0] m. We evaluate both methods based on 6 important metrics as presented in TABLE LABEL:sim_task3_tabular. To offer a more detailed view, Fig. 7 illustrates the quadrotor and payload trajectories of a representative track governed by these two methods.
Figure 7: Time-series comparison of the trajectories for our method versus Impactor [12] on a representative gate traversal track. The shaded regions highlight the traversal phase (blue: our method; orange: Impactor), defined as the duration from the system’s initial entry to its full exit of the gate.
The experimental results indicate that both methods can successfully navigate all 8 trials without collision, and the deviations from the gate center are numerically similar. However, our method exhibits significantly greater agility, attaining a peak velocity 3x higher than Impactor and completing the trajectories in approximately 1/3 of the time. Notably, although Impactor can autonomously switch cable mode (slack or taut), this optimization is computationally demanding. Conversely, our deployed policy navigates the suspended payload system at 100 Hz (matching the training frequency, but flexible in practice), achieving real-time planning and control. This capability enables our approach to handle more complex and dynamic scenarios effectively.
IV EXPERIMENT SETUP AND RESULTS
For the real-world flights, we use a 305 g quadrotor (with a thrust-to-weight ratio of 3.5) equipped with a 70 g payload, as depicted in Fig. 8. The quadrotor runs Betaflight for low-level control and uses a Cool Pi computer for onboard policy inference at 100 Hz via LibTorch. An Optitrack motion capture system provides position and attitude measurements in a $5.0\,\text{m}\times 5.0\,\text{m}\times 2.0\,\text{m}$ space. Furthermore, we directly deploy the simulation-trained network without fine-tuning.
Figure 8: Quadrotor suspended payload system and real-world setup, including a bar target (for scenario 2) and a circular gate (for scenario 3). Notably, the combined size of the system exceeds the gate’s diameter, prohibiting a direct traversal and necessitating agile maneuvers.
IV-1 Agile waypoint passing
We design three distinct tracks to evaluate the performance of the learned policy in scenario 1, as one representative trajectory shown in Fig. 1(a). The 3D trajectories of both the quadrotor and the payload, presented in Fig. 9, demonstrate that the policy can effectively tackle diverse tracks in the waypoint-passing scenario.
Figure 9: The 3D trajectories of both the quadrotor and the payload on 3 real-world tracks in scenario 1.
TABLE V: Performance Comparison of Simulation and Real-World Experiments on 3 Different Tracks
Track Experiment Type Time (s) Max Vel. (m/s)
Track 1
(4 waypoints)
Simulation 1.33 4.71
Real-World 1.24 5.29
Track 2
(7 waypoints)
Simulation 2.36 5.40
Real-World 2.56 5.80
Track 3
(14 waypoints)
Simulation 6.00 4.72
Real-World 6.39 4.86
Additionally, to assess the sim-to-real transfer performance of the learned policy, we also compare the results of the aforementioned tracks as executed in both simulation and the real world. The quantitative results, summarized in Table V, confirm a successful policy transfer, showing that the real-world performance is highly comparable to the simulation with minor deviations.
IV-2 Payload targeting
The performance of the learned policy in scenario 2 is assessed by its ability to reach a given target at a distance of 1.5 m, as shown in Fig. 1(b). The motion data recorded in several trials shows that the policy achieves an average payload targeting error of 0.12 m. The payload’s average and maximum velocities are 1.61 m/s and 3.74 m/s, respectively. The quadrotor’s average and maximum velocities are 1.42 m/s and 2.98 m/s, respectively. These results indicate that the system exhibits high targeting accuracy and excellent speed performance in this scenario.
IV-3 Gate traversal
The system’s gate traversal capability in scenario 3 is assessed by requiring it to navigate through both a single gate and two consecutive gates, as shown in Fig. 1(c). Notably, this scenario is particularly demanding as the combined size of the quadrotor and swinging payload exceeds the gate’s diameter, necessitating agile maneuvers, as the specific configuration detailed in Fig. 8.
In this scenario, both the quadrotor and the payload successfully navigate through narrow gates. During the consecutive traversal, the deviations from the center of the first gate are 0.21 m (quadrotor) and 0.17 m (payload), followed by deviations of 0.25 m and 0.26 m at the second. The entire maneuver is completed in a swift 1.57 s, with the quadrotor achieving an average velocity of 2.19 m/s and a maximum velocity of 5.72 m/s. These results underscore our policy’s remarkable aenility and safety.
V CONCLUSIONS
In this letter, we introduce FLARE, a reinforcement learning framework designed to solve the challenging problem of agile flight for the quadrotor suspended payload system. Our approach, trained entirely in simulation, is validated across three demanding scenarios, demonstrating successful zero-shot sim-to-real transfer. This results in superior agility and computational efficiency compared to traditional optimization-based approaches.
This work highlights that model-free RL is an effective and practical approach for unlocking the vast agility potential of the suspended payload system. Future work could extend this framework to handle dynamic obstacles or more complex interaction tasks.